The distance of station from the house of Ramakant Babu is 8km.He covers this distance in 36 minutes,travelling first part at the speed of 10km. per hour by bi-cycle and the second part at the speed of 18km. per hour by auto-rickshaw.What distance does he travel in each part? : The distance of station from the house of Ramakant Babu is 8km.He covers this distance in 36 minutes,travelling first part at the speed of 10km. per hour by bi-cycle and the second part at the speed of 18km. per hour by auto-rickshaw.What distance does he travel in each part?
Answer by stanbon(18999)

The distance of station from the house of Ramakant Babu is 8km.He covers this distance in 36 minutes,traveling first part at the speed of 10km. per hour by bi-cycle and the second part at the speed of 18km. per hour by auto-rickshaw.What distance does he travel in each part?
-------------
Bicycle DATA:
distance = x km ; rate = 10 km/hr ; time = d/r = x/10 hrs.
---------------
Rickshaw DATA:
distance = (8-x) km ; rate = 18 km/hr ; time = (8-x)/18 hrs
---------------
EQUATION:
time + time = 36/60 hrs
x/10 + (8-x)/18 = 36/60
x/5 + (8-x)/9 = 36/30
Multiply thru by 90 to get
18x + 80-10x = 108
8x = 28
x = 7/2 kmh (bicycle distance)
8-x = 9/2 kmh (rickshaw distance)
===================================
Cheers,
Stan H.